### v1

# Modified Schwarzschild spacetime

Abstract
Considering an ansatz for $1+3+n$ dimensional static spacetime with uniform $n$-branes stacked in three-dimensional rotationally symmetric space, we found different classes of vacuum solutions of Einstein field equations. We analyzed these solutions through the Newtonian limit, Kretschmann scalar, and conserved energy from Landau–Lifshitz stress-energy pseudotensor. This revealed unusual physical properties for a class of solutions with nontrivial extension of the Schwarzschild spacetime with extra dimensions. Positive conserved energy is associated with repulsive gravity and physical singularity on the horizon, while negative conserved energy implies gravitational attraction and naked singularity. This sets the Schwarzschild solution in a broader context and provides new possible testing grounds for general relativity.
1 Introduction
Schwarzschild spacetime [2] is one of the most prominent examples of vacuum solutions of Einstein field equations. Since vacuum solutions correspond to zero Ricci tensor, they can be called Ricci-flat. The Schwarzschild solution can be interpreted as a black hole [3], and such objects are present in our Universe [4, 5, 6]. Black holes may generate Hawking radiation [7], providing a laboratory for testing quantum effects in gravity. Modification of the original Schwarzschild solution to an arbitrary number of space dimensions was first found by Tangherlini [8]. Rotating black holes with Kerr metric [9] can also be generalized to an arbitrary number of space dimensions. They are known as Myers–Perry black holes [10, 11]. There are also other different higher dimensional generalizations [12] including, for example, black rings [13].
Physical motivation for extra dimensions originates mainly from Kaluza–Klein models [14, 15] attempting to solve hierarchy problem [16, 17] and string theory [18, 19, 20] developed as a promising candidate for a quantum theory of gravity. These theories usually require some kind of compactification of extra dimensions, which contradicts the aforementioned examples. Small extra dimensions can be present in a different class of objects known as black strings, p-branes or black branes [21, 22, 23, 24, 25, 26]. They may have black hole properties in macroscopic dimensions while containing compactified extra dimensions from String theory[27, 28, 29, 30]. Higher dimensional objects provide valuable testing grounds for studying quantum effects in general relativity[31, 32, 33].
This paper provides an analysis of a special class of Ricci-flat spacetimes in $1+3+n$ dimensions under the assumption of spherical symmetry in the three-dimensional part and Euclidean symmetry in the $n$-dimensional part, $SO(3)\times E(n)$. Extra dimensions may be compactified into a flat $n$-torus with $S^{1}\times...\times S^{1}$ topology. The number of the standard space dimensions that cannot be compactified is fixed to three in this work since it is the physically most relevant case. Disregarding $1+3+n$ dimensional Minkowski spacetimes, there will be two different classes of solutions. The first class represents a simple extension of the $1+3$ dimensional Schwarzschild spacetime with the addition of $n$ flat dimensions. In contrast, the second class introduces a nontrivial form of $g_{AB}$ components of the spacetime metric, where indices $A$ and $B$ run through $n$ additional dimensions.
The following section 2 provides vacuum solutions of Einstein field equations with two classes mentioned above, section 3 includes further analysis of properties of these spacetimes based on the Newtonian limit, Kretschmann scalar invariant and conserved energy from Landau–Lifshitz pseudotensor, and section 4 is dedicated to summary and discussion. The usual convention with the light speed set to unity, $c=1$, together with the spacetime metric signature (,+,+,)(-,+,+,...)( - , + , + , … ), will be used throughout this paper. The index 00 will be used for a time coordinate, lowercase Latin indices will indicate three space coordinates, $i=1,2,3$, and capital Latin indices will run through $n$ extra dimensions, $A=1,...,n$. All spacetime indices will be collectively denoted by Greek letters, $\mu=0,i,A$.
2 Ricci-flat solutions
The existence of two different classes of Ricci-flat extensions of the Schwarzschild spacetime is obvious in the case with only one extra space dimension, often called black string. The metric of the trivial case can be written as
$ds^{2}=-\left(1+\dfrac{a}{r}\right)dt^{2}+\left(1+\dfrac{a}{r}\right)^{-1}dr^{2}+r^{2}d\Omega_{(2)}^{2}+d\zeta^{2},$ (2.1)
where $t$ is time, $a$ is a constant, $r$ is radial coordinate, $d\Omega_{(2)}^{2}=d\vartheta^{2}+\sin^{2}\vartheta d\varphi^{2}$ is line element of a unit $2$-sphere, and $\zeta$ denotes coordinate in the additional dimension. Since the extra dimension is added most trivially, the spacetime with this metric is Ricci-flat like the original Schwarzschild spacetime. The Ricci tensor remains zero if this metric is rewritten through a formal change $\zeta\to it$, $t\to i\zeta$. The resulting metric reads
$ds^{2}=-dt^{2}+\left(1+\dfrac{a}{r}\right)^{-1}dr^{2}+r^{2}d\Omega_{(2)}^{2}+\left(1+\dfrac{a}{r}\right)d\zeta^{2}.$ (2.2)
Metrics (2.1) and (2.2) describe two spacetimes with different properties, which, for example, can be tested by a slowly moving particle far away from the center in the Newtonian limit. This simple example with only one extra space dimension can be generalized to the arbitrary number of extra dimensions, which will be examined in what follows.
Our goal is to find a metric of a Ricci-flat spacetime with three-dimensional spherical symmetry and $n$-dimensional Euclidean symmetry in extra dimensions, $SO(3)\times E(n)$, which can be interpreted as a stack of uniform $n$-branes in $1+3+n$ dimensions. We start with the ansatz
$ds^{2}=-f^{\alpha}dt^{2}+f^{\beta}dr^{2}+r^{2}d\Omega_{(2)}^{2}+f^{\gamma}\delta_{AB}d\zeta^{A}d\zeta^{B},$ (2.3)
where $\alpha$, $\beta$ and $\gamma$ are constants, and $f=f(r)$ is a function of only the radial coordinate. A more general ansatz may be parameterized by three independent functions of the radial coordinate [24], but in our analysis, we assume that these functions are given by different powers of one function $f(r)$. There are two reasons to start with the special choice of ansatz (2.3). A pragmatic reason is that with this choice, the condition of Ricci tensor being zero reduces to simple algebraic relations. The second reason is that one would expect only one potential event horizon, which occurs only when all components of the spacetime metric $g_{00}$, $g_{rr}$ and $g_{AB}$ are singular at the same radial coordinate, and the simplest way to achieve this is to consider our form of the ansatz. In principle, there may be Ricci-flat solutions with $SO(3)\times E(n)$ symmetry exhibiting more than one horizon; however, with the ansatz (2.3), we will not be able to find them. For example, authors of [34] consider a general form of ansatz with one additional dimension.
All nonzero components of the Ricci tensor in spacetime with the metric of the form (2.3) can be written as
$R_{00}=\alpha f^{\alpha-\beta}F_{1},\quad R_{rr}=F_{2},\quad R_{\vartheta\vartheta}=F_{3},$
$R_{\varphi\varphi}=F_{3}\sin^{2}\vartheta,\quad R_{AB}=-\gamma f^{\gamma-\beta}F_{1}\delta_{AB},$ (2.4)
where
$F_{1}=\dfrac{1}{r}\dfrac{f^{\prime}}{f}+\dfrac{1}{4}\left(\alpha-\beta+n\gamma-2\right)\left(\dfrac{f^{\prime}}{f}\right)^{2}+\dfrac{1}{2}\dfrac{f^{\prime\prime}}{f},$
$F_{2}=\beta\dfrac{1}{r}\dfrac{f^{\prime}}{f}+\dfrac{1}{4}\left[\alpha\left(-\alpha+\beta+2\right)+n\gamma(\beta-\gamma+2)\right]\left(\dfrac{f^{\prime}}{f}\right)^{2}-\dfrac{1}{2}\left(\alpha+n\gamma\right)\dfrac{f^{\prime\prime}}{f},$
$F_{3}=1-f^{-\beta}\left[1+\dfrac{1}{2}\left(\alpha-\beta+n\gamma\right)r\dfrac{f^{\prime}}{f}\right],$ (2.5)
with prime denoting differentiation with respect to the radial coordinate. All three functions $F_{1}$, $F_{2}$, and $F_{3}$ must be zero to have the Ricci-flat solution.
The condition $F_{3}=0$ represents a simple differential equation that can be solved by integration. The result reads
$f=\left(1+ar^{q}\right)^{-1/\beta},\quad q=\dfrac{2\beta}{\alpha-\beta+n\gamma},$ (2.6)
where $a$ is an integration constant, and $q$ is given by constants from the ansatz. This solution is invalid if $\alpha-\beta+n\gamma=0$, and such case has to be treated separately. It is easy to see that in this case, the condition $F_{3}=0$ implies $f=1$, and we retrieve the simple $1+3+n$ dimensional Minkowski spacetime. The same is true if the integration constant $a$ is zero. By inserting (2.6) into expressions for $F_{1}$ and $F_{2}$ in (2.5) we find
$F_{1}=C_{1}\Phi,\quad F_{2}=(C_{2}+C_{3}ar^{q})\Phi,$ (2.7)
where function $\Phi$ and constants $C_{1}$, $C_{2}$ and $C_{3}$ are
$\Phi=\dfrac{1}{\left(\alpha-\beta+n\gamma\right)^{2}}\dfrac{ar^{q-2}}{\left(1+ar^{q}\right)^{2}},$
$C_{1}=-\alpha-\beta-n\gamma,$
$C_{2}=2\beta^{2}+\beta\left(\alpha+n\gamma\right)-\left(\alpha+n\gamma\right)^{2},$
$C_{3}=2\beta^{2}-\alpha^{2}-n\gamma^{2}-\left(\alpha+n\gamma\right)^{2}.$ (2.8)
The problem of finding Ricci-flat solutions of the form (2.3) is now reduced to the system of algebraic equations.
First two conditions $C_{1}=0$ and $C_{2}=0$ are satisfied if $\alpha=-\beta-n\gamma$, which implies $q=-1$, and by inserting this expression for $\alpha$ into the third relation we also find
$C_{3}=-n\gamma\left[2\beta+\left(n+1\right)\gamma\right].$ (2.9)
The condition $C_{3}=0$ is then satisfied by two sets of possible solutions. One set corresponds to the choice $\gamma=0$, and for the second set, we have $02\beta+(n+1)\gamma=$ 0. With the choice $\gamma=0$ we have $\alpha=-\beta$. Components of the spacetime metric are then given by $f^{\alpha}=1+ar^{-1}$, $f^{\beta}=\left(1+ar^{-1}\right)^{-1}$, and $f^{\gamma}=1$, and the corresponding spacetime metric can be written as
$ds^{2}=-\left(1+\dfrac{a}{r}\right)dt^{2}+\left(1+\dfrac{a}{r}\right)^{-1}dr^{2}+r^{2}d\Omega_{(2)}^{2}+\delta_{AB}d\zeta^{A}d\zeta^{B}.$ (2.10)
This is only a trivial extension of the original $1+3$ dimensional Schwarzschild solution, with metric (2.1) as a special case for $n=1$. A more interesting solutions correspond to the second choice with $02\beta+(n+1)\gamma=$ 0, where $\alpha/\beta=(n-1)/(n+1)$, and $\gamma/\beta=-2/(n+1)$. Spacetime metric then can be written as
$ds^{2}=-\left(1+\dfrac{a}{r}\right)^{-\frac{n-1}{n+1}}dt^{2}+\left(1+\dfrac{a}{r}\right)^{-1}dr^{2}+r^{2}d\Omega_{(2)}^{2}+\left(1+\dfrac{a}{r}\right)^{\frac{2}{n+1}}\delta_{AB}d\zeta^{A}d\zeta^{B}.$ (2.11)
We can also check that for $n=1$ we retrieve (2.2). In the case with no extra dimensions, $n=0$, both sets of solutions (2.10) and (2.11) automatically reduce to the Schwarzschild solution. These solutions contain potential singularities. They may also contain horizons because vector field with components $\delta^{\mu}_{0}$ is a Killing vector in all of these spacetimes, and if $g_{00}=0$, the norm of such killing vector is zero. This will be discussed in more detail in the next section.
Different solutions similar to (2.11) can be found in the literature. Authors of [21] worked with low energy string theory action in ten dimensions, and they found a brane solution with a single Ramond-Ramond charge with nontrivial $g_{AB}$ components in terms of our notations. This has been studied further in many other works[35, 36, 37, 38, 39, 40, 41, 42]. Extra dimensions parameterized by coordinates $\zeta^{A}$ in (2.11) may be macroscopic with $\mathbb{R}^{n}$ topology, microscopic with $n$-torus topology, or a combination of both cases with some macroscopic and some microscopic dimensions. In all cases, there is Euclidean symmetry in extra dimensions, but one can assign extra dimensions to the part that is only spherically symmetric [43]. Note also that considering extra time dimensions instead of extra space dimensions, or having mixed time and space extra dimensions, would not change the results because changing signs of $g_{AB}$ components of the metric keeps Ricci tensor being zero. Compactification of the time dimension may also be considered, with implications for branes [44]. In the following section, where we will study some properties of the found solutions (2.10) and (2.11), we will assume the physically most relevant case with three macroscopic space dimensions with coordinates $x^{i}$ and $n$ compactified extra space dimensions with coordinates $\zeta^{A}$.
3 Physical properties
The Newtonian limit provides a very important insight into properties of branes containing spherically symmetric $1+3$ dimensional part [45, 46]. We will examine solutions (2.10) and (2.11) in this limit. To focus on physically most relevant cases, let us assume that all extra dimensions are microscopic with $S^{1}\times...\times S^{1}$ topology of a flat $n$-torus. Space is then $L$-periodic in $n$ space directions, $\zeta^{A}\in\left[0,L\right)$. Macroscopic objects with a size larger than $L$ then can move only in the $1+3$ dimensional section. The same is true for sufficiently long wavelength photons or particles with large enough de Broglie wavelengths. From the perspective of such objects, the spacetime in regions far from the origin, $r\gg|a|$, can be approximated as
$ds^{2}\approx-(1+2\phi)dt^{2}+(1-2\psi)\delta_{ij}dx^{i}dx^{j},$ (3.1)
where $\phi$ is the Newtonian potential. Without any extra microscopic dimensions, it can be shown that $\psi=\phi$ in harmonic coordinates. The geodesic equation for a slowly moving object reduces to the Newtonian equation of motion, $\ddot{x}^{i}\approx-\phi_{,i}$, in this limit. In the case with trivial extension of the Schwarzschild spacetime (2.10), the form of the Newtonian potential is usual, $\phi=(a/2)r^{-1}$, and in the nontrivial case (2.11) we have
$\phi=-\dfrac{n-1}{n+1}\dfrac{a}{2}\dfrac{1}{r}=-\kappa M\dfrac{1}{r},$ (3.2)
where $M$ is the mass of a gravitating body in the context of classical mechanics, which is valid in the Newtonian limit. This provides the standard way of fixing the constant $a$ in the Schwarzschild solution as $a=-2\kappa M$. A consequence of a negative value of $a$ is the existence of a horizon. However, in the modified case with the spacetime metric (2.11), the sign is opposite, which implies naked singularities for positive masses $M$. Negative $a$ and the existence of horizon is then associated with negative mass $M$, which means repulsive gravity in the Newtonian limit. Other examples of different types of black holes with such property are studied, for example, in [47]. We may also notice that the presence of extra dimensions does not change the position of the horizon or its properties considerably. For example, a contrary case can be found in [48]. Note also that for $n=1$, the spacetime metric in the nontrivial case is (2.2), and the Newtonian potential vanishes, which means zero mass $M$.
In the case of the ordinary Schwarzschild metric, there is no physical singularity on the horizon. This can be easily revealed by the calculation of the Kretschmann scalar invariant defined through Riemann curvature tensor as $K=R^{\mu\nu\lambda\sigma}R_{\mu\nu\lambda\sigma}$. For the Schwarzschild spacetime it is $K_{\textrm{Schw.}}=12a^{2}r^{-6}$, indicating singularity only at the origin, and the same is true also for the trivial extension (2.10). In the nontrivial case (2.11) we have
$K=\dfrac{12a^{2}}{r^{6}}\left[1-\dfrac{1}{3}\dfrac{n(n-1)}{(n+1)^{3}}\left(1+\dfrac{r}{a}\right)^{-2}\left(\dfrac{3}{4}n+1+(n+1)\dfrac{r}{a}\right)\right].$ (3.3)
We can see that for $n=0$ and $n=1$, there is no difference from the Schwarzschild spacetime, $K=K_{\textrm{Schw.}}$.
To analyze the behavior of the Kretschmann scalar, it is convenient to rearrange the relation (3.3) into the form
$\dfrac{1}{12}r^{4}K=\dfrac{1}{\sigma^{2}}\left[1-\dfrac{1}{3}\dfrac{n(n-1)}{(n+1)^{3}}\dfrac{1}{\left(1+\sigma\right)^{2}}\left(\dfrac{3}{4}n+1+(n+1)\sigma\right)\right],$ (3.4)
where $\sigma=r/a$. With the use of the dimensionless radial coordinate $\sigma$, we can depict the behavior of the Kretschmann scalar in two cases with different signs of constant $a$ in one plot. Positive values of $\sigma$ correspond to positive constant $a$, and negative $\sigma$ corresponds to negative $a$ because the radial coordinate $r$ may be only positive. The right-hand side of (3.4) is plotted in Fig. 1 for some numbers of extra dimensions $n$, including the limit $n\to\infty$.
Fig. 1: Right-hand side of (3.4) encoding behaviour of the Kretschmann scalar (3.3) as a function of dimensionless radial coordinate $\sigma$ for various numbers of extra space dimensions.
Unless $a=0$, there is always the central singularity at $r=0$. In the case with positive $a$, there is no additional singularity, but for negative values of constant $a$, the Kretschmann scalar indicates a singularity at the horizon, $r=-a$ or $\sigma=-1$. Other examples of such singularities can be found, for example, in [49, 50].
Another hint of a horizon singularity for negative values of $a$ can be found by investigating radial null geodesics. For an ingoing null geodesic, we can use a parameterization of the radial coordinate as $r=-a+\Delta-\lambda$, where $\Delta$ is a positive constant, and $\lambda$ is the geodesic parameter with values in the interval $\lambda\in[0,\Delta]$. It is sufficient to assume the smallness of $\Delta$ compared to $|a|$. At $\lambda=0$ the ingoing signal starts near the horizon, and at $\lambda=\Delta$ it reaches it. In the case with the trivial extension of the Schwarzschild spacetime (2.10), the amount of the coordinate time needed for a signal to complete this path can be calculated as
[…]
and this integral diverges. This indicates only a coordinate type of singularity because the derivative of the radial coordinate with respect to the parameter of the null geodesic is regular on the horizon, and the same is true also for mass particles radially approaching the horizon, where the derivative is with respect to their proper time. This is the same behavior as with the Schwarzschild metric without extra dimensions. In the nontrivial case with the spacetime metric (2.11) for the considered ingoing signal, the amount of coordinate time needed to reach the horizon is given by
[…]
where the last expression is the result in the leading order of the approximation of small $\Delta/|a|$ for $n>0$. We can see that this result is finite which indicates that singularity of the $g_{00}$ component of the metric is not caused by an inappropriate choice of the time coordinate, but rather it suggests a physical singularity, already revealed by the Kretschmann scalar.
Another approach to characterization of properties of spacetimes is based on Landau–Lifshitz stress-energy pseudotensor [51] providing definition of conserved quantities. In a $D$-dimensional spacetime a conserved $D$-momentum $P^{\mu}$ associated with a $D-1$ dimensional space-volume region $\Omega$ can be calculated as
[…]
where $d\Sigma_{\nu}$ denotes $D-2$ dimensional oriented area element of the boundary of $\Omega$, and through object $h^{\mu\nu\lambda}$ we can obtain the total stress-energy pseudotensor as $t_{\textrm{tot}}^{\mu\nu}=h^{\mu\nu\lambda}_{\phantom{\mu\nu\lambda},\lambda}$. It is defined as
$h^{\mu\nu\lambda}=\dfrac{1}{16\pi\kappa}\left[(-g)\left(g^{\mu\nu}g^{\lambda\sigma}-g^{\mu\lambda}g^{\nu\sigma}\right)\right]_{,\sigma}.$ (3.8)
For evaluation of the conserved quantities defined above for spacetimes with metrics (2.10) and (2.11), we chose the $3+n$ dimensional volume $\Omega$ as a sphere in the three-dimensional space part, so that the topology of its boundary $\partial\Omega$ is $S^{2}\times S^{1}\times...\times S^{1}$. The area element is given by $d\Sigma_{i}=dS_{i}$ and $d\Sigma_{A}=0$, where $dS_{i}$ corresponds to the area element of the sphere boundary. The only nonzero component of the $(4+n)$-momentum is energy $\mathcal{E}=P^{0}$, and it can be calculated as
[…]
To evaluate this integral, it is useful to replace the radial coordinate $r$ with coordinate $\rho$ defined by the transformation $r=\rho\left(1-a/(4\rho)\right)^{2}$. This new coordinate is usually called isotropic radial coordinate because in the case with $1+3$ dimensional Schwarzschild metric, the space part of the metric becomes conformally Euclidean. In our case, this is true only for the three-dimensional and $n$-dimensional sections separately. For positive values of $a$, the isotropic radial coordinate $\rho$ cowers the entire spacetime except for the central singularity twice, while for negative $a$, it cowers twice only the region above the horizon. To avoid the double coverage, we may consider $\rho$ only in the interval $[1/4,\infty)$.
In the trivial case (2.10) we have
$ds^{2}=-\left(\dfrac{\chi_{+}}{\chi_{-}}\right)^{2}dt^{2}+\left(\chi_{-}\right)^{4}\delta_{ij}dx^{i}dx^{j}+\delta_{AB}d\zeta^{A}d\zeta^{B},$ (3.10)
where functions $\chi_{+}$ and $\chi_{-}$ are defined as $\chi_{\pm}=1\pm a/(4\rho)$, and space coordinates $x^{i}$ are Cartesian such that $(x^{1})^{2}+(x^{2})^{2}+(x^{3})^{2}=\rho^{2}$. Again, this is only a trivial extension of the Schwarzschild solution, now written in isotropic coordinates. The conserved energy calculated through formula (3.9) for boundary $\partial\Omega$ chosen in the aforementioned way, at some fixed value of the radial coordinate $\rho$, is
$\mathcal{E}(\rho)=-\dfrac{a}{2\kappa}\left(\chi_{-}\right)^{7},$ (3.11)
and in the limit of $\Omega$ covering entire spacetime, $\rho\to\infty$, we have $\mathcal{E}=-a/(2\kappa)=M$. This is the same result as in the case with the Schwarzschild spacetime without additional space dimensions.
In the nontrivial case (2.11), the spacetime metric can be rewritten as
$ds^{2}=-\left(\dfrac{\chi_{+}}{\chi_{-}}\right)^{-2\frac{n-1}{n+1}}dt^{2}+\left(\chi_{-}\right)^{4}\delta_{ij}dx^{i}dx^{j}+\left(\dfrac{\chi_{+}}{\chi_{-}}\right)^{\frac{2}{n+1}}\delta_{AB}d\zeta^{A}d\zeta^{B}.$ (3.12)
The formula (3.9) then implies
$\mathcal{E}(\rho)=\dfrac{a\left(\chi_{+}\right)^{\frac{4n}{n+1}}\left(\chi_{-}\right)^{4\frac{n+2}{n+1}}}{4(n+1)\kappa}\left(\dfrac{n}{\chi_{+}}-\dfrac{n+2}{\chi_{-}}\right),$ (3.13)
and for the entire spacetime we have $\mathcal{E}=-a/(2(n+1)\kappa)$. We may also notice that for $n=0$, we retrieve the result of the trivial case (3.11).
In summary, we have $a=-2\kappa M$, and $\mathcal{E}=M$ for the Schwarzschild spacetime as well as for its trivial extensions (2.10). On the other hand, in the nontrivial case
$a=2\dfrac{n+1}{n-1}\kappa M,\quad\mathcal{E}=-\dfrac{1}{n-1}M,$ (3.14)
and there is also horizon singularity at $r=-a$ if $a$ is negative and $n>1$. For $n=0$, both cases coincide with the Schwarzschild spacetime, and in the nontrivial case with $n=1$, the mass given by the Newtonian limit is zero, $M=0$, and constant $a$ cannot be fixed in the usual way. These unusual properties set the nontrivial case (2.11) apart from the trivial one (2.10).
4 Discussion
There are three distinctive classes of Ricci-flat spacetimes with metric given by the ansatz (2.3). For $f=1$, we have $1+3+n$ dimensional Minkowski spacetime. We focused on more intriguing cases given by metrics (2.10) and (2.11). The first is a direct product of the $1+3$ dimensional Schwarzschild spacetime and $n$-dimensional Euclidean space. If we assume microscopic extra dimensions with $n$-torus topology, such spacetimes do not differ from the ordinary Schwarzschild spacetime from the point of view of physical objects larger than the size of extra dimensions. Most interesting Ricci-flat solutions are given by the metric (2.11), and in the previous section, we found their nonstandard properties.
Comparing them with other cases mentioned above is helpful in appreciating their unusual features. This can be summarized in table 1.
horizons singularities Newt. lim. mass cons. energy 1. Minkowski ($a=0$) none none $M=0$ $\mathcal{E}=0$ 2. trivial ($a>0$) none $r=0$ $M=\frac{-a}{2\kappa}<0$ $\mathcal{E}=M<0$ 3. trivial ($a<0$) $r=-a$ $r=0$ $M=\frac{-a}{2\kappa}>0$ $\mathcal{E}=M>0$ 4. nontrivial ($a>0$) none $r=0$ $M=\frac{n-1}{n+1}\frac{a}{2\kappa}>0$ $\mathcal{E}=\frac{-a}{2(n+1)\kappa}<0$ 5. nontrivial ($a<0$) $r=-a$ $r=\{0,-a\}$ $M=\frac{n-1}{n+1}\frac{a}{2\kappa}<0$ $\mathcal{E}=\frac{-a}{2(n+1)\kappa}>0$
Table 1: Summary of properties of all possible Ricci-flat spacetimes with the metric of the form (2.3), including Minkowski spacetime and solutions (2.10) and (2.11). Note that in case 4., the existence of the horizon singularity at $r=-a$ occurs only for the number of extra dimensions $n>1$.
As we can see, only cases 1. and 3 are well-behaved and have the same properties as Minkowski and Schwarzschild spacetimes. Case 2. is nonphysical because it requires negative mass defined through the Newtonian limit and conserved energy from the Landau–Lifshitz pseudotensor, which also coincide with each other, $\mathcal{E}=M$. It also lacks a horizon, and therefore, it has a naked singularity. These reasons also exclude the possibility of negative values of constant $a$ for the Schwarzschild solution with no extra dimensions.
Cases 4. and 5. with spacetime metric of the form (2.11) show some peculiar properties. The sign of constant $a$ determines the existence of a horizon. It is located at $r=-a$, if $a$ is negative. There is a central singularity at $r=0$ in both cases, as indicated by the Kretschmann scalar (3.3), but there is also a horizon singularity in the case with negative $a$ and $n>1$. In case 4., there is a naked singularity, attractive gravitational effect in the Newtonian limit, $M>0$; However, the conserved energy $\mathcal{E}$ is negative. In case 5., there is a horizon singularity, the conserved energy is positive, but gravity acts repulsively in the Newtonian limit.
Another important difference between cases 2. and 3. with the spacetime metric (2.10) and cases 4. and 5. with (2.11) is the size of extra dimensions, if they are compactified with the $n$-torus topology. With the trivial extension of the Schwarzschild spacetime, we have $g_{AB}=\delta_{AB}$. Hence, the size of extra dimensions remains the same throughout the entire spacetime. However, we have $g_{AB}=\left(1+a/r\right)^{2/(n+1)}\delta_{AB}$ in the nontrivial case. In the case with positive $a$, there is no horizon, and the size of extra dimensions smoothly transits from a constant in large distances from the center to infinite divergence at the central singularity. This behavior at the center and far away is the same for the case with negative $a$, but in this case, the size of extra dimensions also drops to zero at the horizon. As we have shown, this horizon corresponds to a physical singularity.
Solutions with the nontrivial extension of the Schwarzschild spacetime (2.11) may, in principle, correspond to exterior solutions around some spherically symmetric objects [52]. However, such objects would have nonphysical properties. If we demand the conserved energy $\mathcal{E}$ to be positive, such an object would manifest repulsive gravity. A system containing such an exotic object and another massive astrophysical object could then accelerate itself, which is physically nonsensical. To avoid this problem, the conserved energy $\mathcal{E}$ must be negative, which is also problematic. This suggests that the solution (2.11) is nonphysical.
In addition to problems with nonphysical properties, there are potential problems with instability. Some higher-dimensional extensions of the Schwarzschild spacetime tend to be unstable [53], which also includes brane solutions [54, 55]. Case 4. in the table contains naked singularity, which also hints at potential problems with stability [56, 57].
Finally, let us point out that only case 3. in the table can be maximally extended in the same way as the Schwarzschild spacetime [58]. In cases 2., 4. and 5., either the horizon singularity or lack of the horizon prevents the maximal extension by passing to Kruskal–Szekeres coordinates.
This paper has provided an analysis of possible extensions of the Schwarzschild spacetime with $n$ extra dimension and $SO(3)\times E(n)$ symmetry. We may conclude that only solutions with the spacetime metric (2.10) and choice of negative values for constant $a$ are physically well-behaved. They represent a trivial extension of the $1+3$ dimensional Schwarzschild spacetime. We have also found more intriguing solutions (2.11). Despite their problematic features, they provide an interesting testing ground for general relativity since they represent one of not many examples of vacuum solutions of Einstein field equations. Problems with some of their nonphysical properties may perhaps be avoided by generalizing and modifying them further, but we leave this for further work.
Acknowledgement
The work was supported by grants VEGA 1/0719/23, VEGA 1/0025/23, and Ministry of Education contract No. 0466/2022.